Find the cost per pound of a "house blend" of coffee that is made from 16 lb of Central American coffee that costs $7 per pound and 40 lb of South American coffee that costs $3.50 per pound
Answer by jorel555(1290)

16(7)+40(3.50)=112+140=252
252/(40+16)=$4.5
You now have 56 lbs. of coffee worth $4.50/lb. ☺☺☺☺